# Disentangled Latent Energy-Based Style Translation: An Image-Level Structural MRI Harmonization Framework

### v1

@@ 3.2.1 Problem Formulation
In this work, we reformulate image-level MRI harmonization as a domain-level translation task, which is different from previous studies that treat each MRI scan as a unique domain and concentrate on individual source-to-reference image translation (Liu et al., 2021; Cackowski et al., 2023). In the context of multi-site MRI studies, we consider a single site or scanning setting for MRI acquisition as a unique domain. This is justified by numerous studies demonstrating that MRIs scanned in the same site under the same setting can exhibit intra-site variations that result in different styles within one domain (Deoni et al., 2008; Styner et al., 2002). Consequently, image-level translation using only one reference image may not account for these variations, failing to capture true data distributions of a site/domain. For the two datasets used in this work, we first compute the mean peak signal-to-noise ratio (PSNR) for each domain by averaging the PSNR values obtained from every unique pair of images within that same domain. We then designate the domain with minimal intra-site variations (highest PSNR) as the target domain, with the remaining site(s) as the source domain(s). Overall, we aim to translate MRIs within a source domain $X$ to have the imaging style (i.e., non-biological variation) of MRIs in a target domain $Y$.
@@ 3.2.2 Site-Invariant Image Generation (SIG)
As shown in the left panel of Fig. 1, the SIG is designed to encode MRI scans into a lower-dimensional latent space and subsequently decode reconstructed MRIs from these latent codes. Importantly, this process is trained independently of any site specifications, and thus, does not require any site labels or image pairs. Specifically, the SIG module takes as input a real MRI ${\bf x}\sim\mathbb{P}_{Data}$ and a random noisy image $z$ drawn from a Gaussian distribution $\mathcal{N}(0,1)$, where $\mathbb{P}_{Data}$ denotes the real data distribution. A network $\bm{F}$ learns a deterministic mapping to convert $z$ into a latent code (Pidhorskyi et al., 2020), represented as $\mathcal{W}={F}(z)$. Following this, we further design a stochastic decoding network $\bm{G}$ to generate a synthetic MR image based on the latent code $\mathcal{W}$ and an independent noise $\eta\sim\mathcal{N}(0,1)$, represented as $\tilde{{\bf x}}={G}(\mathcal{W},\eta)$. A network $\bm{E}$ is then trained to encode the synthetic image $\tilde{{\bf x}}$ back into the latent space, resulting in a reconstructed latent code $\tilde{\mathcal{W}}$. This entire process can be symbolized as $\tilde{\mathcal{W}}=E(\tilde{{\bf x}})=E\circ G\circ F(z)$, where the symbol $\circ$ denotes function composition. The encoder $\bm{E}$ also can transform a real image ${\bf x}$ into its corresponding latent code, denoted as $\mathcal{R}=E({\bf x})$. Lastly, a latent discriminator, denoted as $\bm{D}$, receives a latent code and produces a probability value that signifies whether the latent code is derived from a synthetic or real MRI. The networks $\bm{F}$, $\bm{G}$, $\bm{E}$, and $\bm{D}$ are trained using an adversarial approach, enabling the SIG module to generate latent codes that closely resemble those derived from real MRIs. The latent code $\mathcal{R}$, derived from the real image, is also decoded back to a reconstructed image using the decoder $\bm{G}$, aiming to preserve the content (e.g., anatomical structures) of the real MRI.
@@
$\small\mathcal{L}_{lae}^{E,G}=\Delta_{1}(F\lVert E\circ G\circ F)=\mathbb{E}_{z\sim\mathbb{P}_{z}}[\lVert F(z)-E\circ G\circ F(z)\rVert_{2}^{2}]$ (1)
Through this process, the network pair $\bm{E}$ and $\bm{G}$ can be interpreted as a latent autoencoder that encodes and reconstructs the latent code $\mathcal{W}$. This differs from conventional autoencoders, which typically operate in high-dimensional image space (Kingma and Welling, 2013). The generation of images via autoencoding low-dimensional latent codes enhances computational efficiency.
In order to maintain the anatomical structure preservation of the latent autoencoder $\bm{E}$ and $\bm{G}$, a pixel-wise content loss is introduced to computer the difference between the original real image ${\bf x}$ and its reconstructed counterpart $\tilde{{\bf x}}$, as follows:
$\small\mathcal{L}_{pix}^{E,G}=\Delta_{2}({\bf x}\lVert\tilde{{\bf x}})=\mathbb{E}_{x\sim\mathbb{P}_{Data}}[\lVert{\bf x}-G\circ E({\bf x})\rVert_{1}]$ (2)
@@
$\small\mathcal{L}_{adv}^{E,D}=\Phi(D\circ E\circ G\circ F(z))+\Phi(-D\circ E({\bf x}))+R_{reg}$ (3)
$\small\mathcal{L}_{adv}^{F,G}=\Phi(-D\circ E\circ G\circ F(z))$ (4)
where $\Phi$ is the $softplus$ function defined as $\Phi(t)=\log(1+\exp(t))$, $R_{reg}$ is a zero-centered gradient penalty term defined as $\frac{\gamma}{2}\mathbb{E}_{x\sim\mathbb{P}_{Data}}[\lVert\nabla D\circ E({\bf x})\rVert^{2}]$ (Pidhorskyi et al., 2020).
Overall, the final objective functions of SIG is as follows:
$\small{\operatorname{min}_{\{F,G\}}}\,{\operatorname{max}_{\{E,D\}}}\,\mathcal{L}_{adv}^{E,D}+\mathcal{L}_{adv}^{F,G},~{}~{}{\operatorname{min}_{\{E,G\}}}\,\mathcal{L}_{lae}^{E,G}+\mathcal{L}_{pix}^{E,G}$ (5)
@@ 3.2.3 Site-Specific Style Translation (SST)
As depicted in the upper right panel of Fig. 1, the proposed SST consists of (1) an autoencoder that includes the networks $\bm{E}$ and $\bm{G}$ trained in the SIG module, and (2) an energy-based model (EBM) that performs latent-space style translation. Given an MRI ${\bf x}$ from source site $X$ and an MRI ${\bf y}$ from target site $Y$, the encoder $\bm{E}$ initially transforms them into their respective latent codes, denoted as $Z_{x}=E({\bf x})$ and $Z_{y}=E({\bf y})$, respectively. These latent codes are then input into the EBM, which produces the translated latent code $Z_{x\to y}$. This translated latent code is subsequently decoded by $\bm{G}$ to produce the translated MRI. The EBM is also utilized to reverse-translate $Z_{x\to y}$, resulting in a recovered source latent code, denoted as $\tilde{Z}_{x}$. This reconstructed latent code is compared with the original $Z_{x}$ to ensure the latent cycle-consistency, which prevents mode collapse during the unpaired style translation process. The specifics of the EBM are discussed as follows.
@@
$\footnotesize\begin{split}\mathcal{L}_{EBM}&=-(\frac{\partial}{\partial\theta}-L(\theta))\\ &=-(\mathbb{E}_{Z_{y}\sim\mathbb{P}_{Data}}[\frac{\partial}{\partial\theta}\mathcal{E}_{\theta}(Z_{y})]-\mathbb{E}_{\tilde{Z}_{y}\sim\mathbb{P}_{\theta}}[\frac{\partial}{\partial\theta}\mathcal{E}_{\theta}(\tilde{Z}_{y})])\end{split}$ (6)
where the first expectation term $\mathbb{E}_{Z_{y}}$ corresponds to samples $Z_{y}$, which are drawn from the actual target data distribution $\mathbb{P}_{Data}(Y)$. On the other hand, the second expectation term $\mathbb{E}_{\tilde{Z}_{y}}$ corresponds to samples $\tilde{Z}_{y}$, which are drawn from the modeled distribution $\mathbb{P}_{\theta}$. This requires direct sampling from $\mathbb{P}_{\theta}$, an intractable model distribution. This can be approximated using a Markov chain Monte Carlo (MCMC) sampling method, namely the Stochastic Gradient Langevin Dynamics (SGLD) (Welling and Teh, 2011), an iterative method defined as follows:
@@
where $t=[0,\cdots,T]$ represents the sampling steps with a step size $\eta$ and $\epsilon$ is a Gaussian noise that is injected to account for data uncertainty and to improve sample convergence (LeCun et al., 2006). The initial $\tilde{Z}_{y}^{0}$ are latent codes sampled from the source site $X$, denoted as $\tilde{Z}_{y}^{0}=Z_{x}=E({\bf x})$. The SGLD sampling process involves iteratively updating the initial source latent codes using negative gradients of the energy function $\mathcal{E}_{\theta}$. This sampling process can be interpreted as evolving original source latent codes towards a configuration with lower energy, thereby aligning them more closely with the true target latent distribution. In contrast to source-to-reference image translation (Liu et al., 2021; Zhu et al., 2017; Choi et al., 2020), EBM effectively traverses the latent space of the entire target domain, allowing a comprehensive consideration for intra-site variations within the target domain. After $T$ sampling steps (via SGLD), we can decode $\tilde{Z}_{y}^{T}$ through $\bm{G}$ to generate a translated MRI, denoted as $\tilde{{\bf y}}=G(Z_{x\to y})=G(\tilde{Z}_{y}^{T})$.
Considering that the EBM does not differentiate the content code and the style code explicitly during style translation, we further introduce a latent content loss into the EBM for regularization. This loss function measures the difference between the translated latent code $Z_{x\to y}$ and the original source latent code $Z_{x}$, which is formulated as:
@@
where $\alpha$ and $\beta$ serve as hyperparameters that determine the relative contributions of the three terms in the equation. By disentangling image generation via SIG and style translation via SST in the low-dimensional latent space, it is expected to achieve efficient and generalizable MRI harmonization.
@@ 3.2.4 Site-Specific MRI Synthesis (SMS)
In the proposed SST module, we can substitute the original source MRI and encoder $\bm{E}$ with a random Gaussian noise $z$ and the mapping network $\bm{F}$ to perform site-specific MRI synthesis. In this case, as the network $\bm{F}$ has already been trained in the SIG to generate latent codes resembling brain MRIs, no further fine-tuning is required. The EBM will then translate these randomly generated latent codes into the target style for image synthesis. Unlike many MRI synthesis models that only generate MRI with random anatomical structures and styles, our framework can synthesize MRIs with a consistent intensity distribution that matches a given image domain, this can be effectively used as a data augmentation method to pre-train models that require large number of MRI samples without introducing non-biological variations caused by difference in imaging style.
Figure 2: Histogram comparison of 10 source sites and a target site (COI) from the SRPBS dataset. The first plot shows pre-harmonization histograms, while the subsequent plots depict post-harmonization histograms by each competing method and our method DLEST, respectively. WM: white matter; GM: gray matter; CSF: cerebrospinal fluid.
@@ 3.2.5 Implementation
The implementation of SIG incorporates similar settings as outlined in (Pidhorskyi et al., 2020), with structural modifications made to accommodate our specific datasets and tasks. Specifically, both the mapping function $\bm{F}$ and the latent discriminator $\bm{D}$ are implemented using multi-layer perceptions (MLPs). The former consists of eight layers, while the latter is composed of three layers. Two networks $\bm{E}$ and $\bm{G}$ in the latent autoencoder are composed of seven encoding/decoding blocks. Each of these blocks incorporates an increasing number of filters, ranging from 64 to 256 filters. All latent codes have a dimension of 512 in the experiments. We train the SIG module using an Adam (Kingma and Ba, 2014) optimizer in an unsupervised training scheme. Compared to GANs (Zhu et al., 2017; Choi et al., 2020), SIG’s training cost is lower due to two factors: (1) most of its losses are evaluated in low-dimensional latent space, and (2) no additional cost for cycle-consistency constraints in image space. Following the SIG training, we can directly use the latent autoencoder in SST training, during which $\bm{E}$ and $\bm{G}$ are not updated.
The training of the SST is efficient in both training time and computational cost due to four factors. (1) The autoencoder $\bm{E}$ and $\bm{G}$ do not require additional training, even when deployed on previously unseen data. (2) The style translation process, facilitated by EBM, operates in a low-dimensional latent space with a dimension of 512, which is considerably smaller than the image space. (3) The energy function $\mathcal{E}_{\theta}$ in EBM is designed to be light-weight, also implemented as an MLP with 2 layers, each having a dimension of 2,048. (4) The SST can be trained with a small sample size for each target site (i.e., $\sim$120120120120 slices in our experiments). In this way, our DLEST can achieve efficient imaging style translation and generalizable MRI generation through the disentanglement scheme.
@@ 4.1.1 Competing Methods
We first compare DLEST with two non-learning methods, including (1) histogram matching (HM(Shinohara et al., 2014) with multiple experiments with randomly selected reference MRI from the target dataset, and (2) Spectrum Swapping-based Image-level Harmonization (SSIMH), which replaces certain low-frequency region of the source MRI with the corresponding low-frequency spectrum of the reference MRI, with a default swapping threshold 3. We use the DomainATM toolbox (Guan and Liu, 2023) to carry out two non-learning methods. We also compare DLEST against the following state-of-the-art (SOTA) methods for unpaired image-level harmonization.
(1) CycleGAN (Zhu et al., 2017) is initially introduced for natural image-to-image translation. It enhances the capabilities of traditional GANs by integrating an important cycle-consistency constraint. This additional constraint enables the unpaired mapping from a source domain to a target domain. The CycleGAN has been utilized in several MRI harmonization studies (Modanwal et al., 2020; Chang et al., 2022) and has become the foundational structure of many subsequent image-level harmonization methods.
(2) StarGAN (Choi et al., 2020) is initially introduced for multi-domain natural image synthesis task. Then, it was employed for MRI harmonization (Liu et al., 2021). Building upon CycleGAN, the StarGAN method learns a separate latent style code for each target image and injects this code into the source image via instance normalization.
(3) ImUnity (Cackowski et al., 2023) is designed specifically for MRI harmonization. It employs a self-supervised variational autoencoder model, coupled with a confusion module for site/scanner-bias unlearning. In addition, it includes an optional biological preservation module to preserve clinical information, such as disease labels or diagnostic scores. In this work, we use the default form of ImUnity without this optional module, since the two datasets only consist of healthy control subjects.
In all three deep learning methods mentioned above, we typically adhere to the default setting of each competing method and make a dedicated effort to ensure other training hyperparameters align closely with the proposed DLEST framework. In DLEST, the two hyperparameters $\alpha$ and $\beta$ in Eq. 10 are empirically set as 100 and 10, respectively. For a fair comparison, we standardize the input size to be $256\times 256$ and set the latent dimension to 512, mirroring the specifications of the proposed DLEST. To train deep learning models (i.e., CycleGAN, StarGAN, ImUnity, and our DLEST), we use the OpenBHB dataset, due to its relatively larger sample size and abundant sites/settings. Following (Dufumier et al., 2022), 3,227 training scans and 757 validation scans from 58 sites/settings are used with manual inspection and quality control.

### v2

@@ 3.2.1 Problem Formulation
In this work, we reformulate image-level MRI harmonization as a domain-level translation task, which is different from previous studies that treat each MRI scan as a unique domain and concentrate on individual source-to-reference image translation [23, 3]. In the context of multi-site MRI studies, we consider a single site or scanning setting for MRI acquisition as a unique domain. This is justified by numerous studies demonstrating that MRIs scanned in the same site under the same setting can exhibit intra-site variations that result in different styles within one domain [6, 37]. Consequently, image-level translation using only one reference image may not account for these variations, failing to capture true data distributions of a site/domain. In this work, we propose to employ all MRIs from a target site as reference images, aiming to better encapsulate the stylistic features of target data. Overall, we aim to translate MRIs within a source domain $X$ to have the imaging style (i.e., non-biological variation) of MRIs in a target domain $Y$.
@@ 3.2.2 Site-Invariant Image Generation (SIG)
As shown in Fig. 1 (a), the SIG is designed to encode MRI scans into a lower-dimensional latent space and subsequently decode reconstructed MRIs from these latent codes. Importantly, this process is trained on the OpenBHB training set, independently of any site specifications, and thus, does not require any site labels or paired images. Specifically, the SIG module takes as input a real MRI ${\bf x}\sim\mathbb{P}_{Data}$ and a random noisy image $z$ drawn from a Gaussian distribution $\mathcal{N}(0,1)$, where $\mathbb{P}_{Data}$ denotes the real data distribution. A network $\bm{F}$ learns a deterministic mapping to convert $z$ into a latent code [29], represented as $\mathcal{W}={F}(z)$. Following this, we further design a stochastic decoding network $\bm{G}$ to generate a synthetic MR image based on the latent code $\mathcal{W}$ and an independent noise $\eta\sim\mathcal{N}(0,1)$, represented as $\tilde{{\bf x}}={G}(\mathcal{W},\eta)$. A network $\bm{E}$ is then trained to encode the synthetic image $\tilde{{\bf x}}$ back into the latent space, resulting in a reconstructed latent code $\tilde{\mathcal{W}}$. This entire process can be symbolized as $\tilde{\mathcal{W}}=E(\tilde{{\bf x}})=E\circ G\circ F(z)$, where the symbol $\circ$ denotes function composition. The encoder $\bm{E}$ also can transform a real image ${\bf x}$ into its corresponding latent code, denoted as $\mathcal{R}=E({\bf x})$. Lastly, a latent discriminator, denoted as $\bm{D}$, receives a latent code and produces a probability value that signifies whether the latent code is derived from a synthetic or real MRI. The networks $\bm{F}$, $\bm{G}$, $\bm{E}$, and $\bm{D}$ are trained using an adversarial approach, enabling the SIG module to generate latent codes that closely resemble those derived from real MRIs. The latent code $\mathcal{R}$, derived from the real image, is also decoded back to a reconstructed image using the decoder $\bm{G}$, aiming to preserve the content (e.g., anatomical structures) of the real MRI.
@@
$\mathcal{L}_{lae}^{E,G}=\Delta_{1}(F\lVert E\circ G\circ F)=\mathbb{E}_{z\sim\mathbb{P}_{z}}[\lVert F(z)-E\circ G\circ F(z)\rVert_{2}^{2}]$ (1)
Through this process, the network pair $\bm{E}$ and $\bm{G}$ can be interpreted as a latent autoencoder that encodes and reconstructs the latent code $\mathcal{W}$. This differs from conventional autoencoders, which typically operate in high-dimensional image space [20]. The generation of images via autoencoding low-dimensional latent codes enhances computational efficiency.
To maintain the anatomical structure preservation of the latent autoencoder $\bm{E}$ and $\bm{G}$, a pixel-wise content loss is introduced to compute the difference between the original real image ${\bf x}$ and its reconstructed counterpart $\tilde{{\bf x}}$, as follows:
$\mathcal{L}_{pix}^{E,G}=\Delta_{2}({\bf x}\lVert\tilde{{\bf x}})=\mathbb{E}_{x\sim\mathbb{P}_{Data}}[\lVert{\bf x}-G\circ E({\bf x})\rVert_{1}]$ (2)
@@
$\mathcal{L}_{adv}^{E,D}=\Phi(D\circ E\circ G\circ F(z))+\Phi(-D\circ E({\bf x}))+R_{reg}$ (3)
$\mathcal{L}_{adv}^{F,G}=\Phi(-D\circ E\circ G\circ F(z))$ (4)
where $\Phi$ is the $softplus$ function defined as $\Phi(t)=\log(1+\exp(t))$, $R_{reg}$ is a zero-centered gradient penalty term defined as $\frac{\gamma}{2}\mathbb{E}_{x\sim\mathbb{P}_{Data}}[\lVert\nabla D\circ E({\bf x})\rVert^{2}]$ [29].
Overall, the final objective functions of SIG are as follows:
${\operatorname{min}_{\{F,G\}}}\,{\operatorname{max}_{\{E,D\}}}\,\mathcal{L}_{adv}^{E,D}+\mathcal{L}_{adv}^{F,G},~{}~{}{\operatorname{min}_{\{E,G\}}}\,\mathcal{L}_{lae}^{E,G}+\mathcal{L}_{pix}^{E,G}$ (5)
@@ 3.2.3 Site-Specific Style Translation (SST)
As depicted in Fig. 1 (b), the proposed SST consists of (1) an autoencoder that includes the networks $\bm{E}$ and $\bm{G}$ trained in the SIG module, and (2) an energy-based model (EBM) that performs latent-space style translation. Given an MRI ${\bf x}$ from source site $X$ and an MRI ${\bf y}$ from target site $Y$, the encoder $\bm{E}$ initially transforms them into their respective latent codes, denoted as $Z_{x}=E({\bf x})$ and $Z_{y}=E({\bf y})$, respectively. These latent codes are then input into the EBM, which produces the translated latent code $Z_{x\to y}$. This translated latent code is subsequently decoded by $\bm{G}$ to produce the translated MRI. The EBM is also utilized to reverse-translate $Z_{x\to y}$, resulting in a recovered source latent code, denoted as $\tilde{Z}_{x}$. This reconstructed latent code is compared with the original $Z_{x}$ to ensure the latent cycle-consistency, which prevents mode collapse during the unpaired style translation process. The specifics of the EBM are discussed as follows.
@@
$\begin{split}\nabla_{\theta}\mathcal{L}_{EBM}&=-\frac{\partial}{\partial\theta}L(\theta)\\ &=\mathbb{E}_{Z_{y}\sim\mathbb{P}_{Data}}\left[\frac{\partial}{\partial\theta}\mathcal{E}_{\theta}(Z_{y})\right]-\mathbb{E}_{\tilde{Z}_{y}\sim\mathbb{P}_{\theta}}\left[\frac{\partial}{\partial\theta}\mathcal{E}_{\theta}(\tilde{Z}_{y})\right]\end{split}$ (6)
where the first expectation term $\mathbb{E}_{Z_{y}}$ corresponds to samples $Z_{y}$, which are drawn from the actual target data distribution $\mathbb{P}_{Data}(Y)$. On the other hand, the second expectation term $\mathbb{E}_{\tilde{Z}_{y}}$ corresponds to samples $\tilde{Z}_{y}$, which are drawn from the modeled distribution $\mathbb{P}_{\theta}$. This requires direct sampling from $\mathbb{P}_{\theta}$, an intractable model distribution. This can be approximated using a Markov chain Monte Carlo (MCMC) sampling method, namely the Stochastic Gradient Langevin Dynamics (SGLD) [44], an iterative method defined as follows:
@@
where $t=[0,\cdots,T]$ represents the sampling steps with a step size $\eta$ and $\epsilon$ is a Gaussian noise that is injected to account for data uncertainty and to improve sample convergence [21]. The initial $\tilde{Z}_{y}^{0}$ are latent codes sampled from the source site $X$, denoted as $\tilde{Z}_{y}^{0}=Z_{x}=E({\bf x})$. The SGLD sampling process involves iteratively updating the initial source latent codes using negative gradients of the energy function $\mathcal{E}_{\theta}$. This sampling process can be interpreted as evolving original source latent codes towards a configuration with lower energy, thereby aligning them more closely with the true target latent distribution. In contrast to source-to-reference image translation [23, 48, 5], EBM effectively traverses the latent space of the entire target domain, allowing a comprehensive consideration for intra-site variations within the target domain. After $T$ sampling steps (via SGLD), we can decode $\tilde{Z}_{y}^{T}$ through $\bm{G}$ to generate a translated MRI, denoted as $\tilde{{\bf y}}=G(Z_{x\to y})=G(\tilde{Z}_{y}^{T})$.
Considering that the EBM does not differentiate the content code and the style code explicitly during style translation, we further introduce a latent content loss into the EBM for regularization measure. This approach is designed to prompt the model to minimally alter the latent code during style translation, thereby preserving the overall content. This loss function measures the difference between the translated latent code $Z_{x\to y}$ and the original source latent code $Z_{x}$, which is formulated as:
@@
Instead of using the latent content loss in Eq. (8), one can also explicitly compute the content loss and the style loss during the latent style translation, which will be discussed in Section 5.3.
@@
where $\alpha$ and $\beta$ serve as hyperparameters that determine the relative contributions of the three terms in the equation. The SST can be trained or fine-tuned for any target style translation with a small sample size for each target site (i.e., $\sim$100100100100 MRI slices in our experiments). The fine-tuning process on a new target site proves to be relatively efficient in terms of time and computation resources. In our experiments, it takes approximately half an hour and utilizes 10G (out of 24G) of GPU memory on an RTX 3090 GPU. More details are discussed in Section 5.5. By disentangling image generation via SIG and style translation via SST in the low-dimensional latent space, it is expected to achieve efficient and generalizable MRI harmonization.
@@ 3.2.4 Site-Specific MRI Synthesis (SMS)
In the proposed SST module, we can substitute the original source MRI and encoder $\bm{E}$ with a random Gaussian noise $z$ and the mapping network $\bm{F}$ to perform site-specific MRI synthesis, as shown in Fig. 1 (c). In this case, as the network $\bm{F}$ has already been trained in the SIG to generate latent codes resembling brain MRIs, no further fine-tuning is required. The EBM will then translate these randomly generated latent codes into the target style for image synthesis. Unlike many MRI synthesis models that only generate MRI with random anatomical structures and styles, our framework can synthesize MRIs with a consistent intensity distribution that matches a given image domain, this can be effectively used as a data augmentation method to pre-train models that require large number of MRI samples without introducing non-biological variations caused by difference in imaging style.
Figure 2: Histogram comparison of 10 source sites and a target site (COI) across all 9 traveling subjects from the SRPBS dataset. The first plot shows pre-harmonization histograms, while the subsequent plots depict post-harmonization histograms by each competing method and our method DLEST, respectively. WM: white matter; GM: gray matter; CSF: cerebrospinal fluid.
@@ 3.2.5 Implementation
The implementation of SIG incorporates similar settings as outlined in [29], with structural modifications made to accommodate our specific datasets and tasks. Specifically, both the mapping function $\bm{F}$ and the latent discriminator $\bm{D}$ are implemented using multi-layer perceptrons (MLPs). The former consists of eight layers, while the latter is composed of three layers. Two networks $\bm{E}$ and $\bm{G}$ in the latent autoencoder are composed of seven encoding/decoding blocks. Each of these blocks incorporates an increasing number of filters, ranging from 64 to 256 filters. All latent codes have a dimension of 512 in the experiments. We train the SIG module using an Adam [19] optimizer in an unsupervised training scheme. Compared to GANs [48, 5], SIG’s training cost is lower due to two factors: (1) most of its losses are evaluated in low-dimensional latent space, and (2) no additional cost for cycle-consistency constraints in image space. Following the SIG training, we can directly use the latent autoencoder in SST training, during which $\bm{E}$ and $\bm{G}$ are not updated.
The training of the SST is efficient in both training time and computational cost due to four factors. (1) The autoencoder $\bm{E}$ and $\bm{G}$ do not require additional training, even when deployed on previously unseen data. (2) The style translation process, facilitated by EBM, operates in a low-dimensional latent space with a dimension of 512, which is considerably smaller than the image space. (3) The energy function $\mathcal{E}_{\theta}$ in EBM is designed to be light-weight, also implemented as an MLP with 2 layers, each having a dimension of 2,048. (4) The SST can be fine-tuned with a small sample size at low computational cost, which will be discussed in Sections 5.4 and  5.5. In this way, our DLEST can achieve efficient imaging style translation and generalizable MRI generation through the disentanglement scheme.
@@ 4.1.1 Competing Methods
We first compare DLEST with two non-learning methods, including (1) histogram matching (HM[34] with multiple experiments with randomly selected reference MRI from the target dataset, and (2) Spectrum Swapping-based Image-level Harmonization (SSIMH), which replaces certain low-frequency region of the source MRI with the corresponding low-frequency spectrum of the reference MRI, with a default swapping threshold 3. We use the DomainATM toolbox [13] to carry out two non-learning methods. We also compare DLEST against the following state-of-the-art (SOTA) methods for unpaired image-level harmonization.
(1) CycleGAN [48] is initially introduced for natural image-to-image translation. It enhances the capabilities of traditional GANs by integrating an important cycle-consistency constraint. This additional constraint enables the unpaired mapping from a source domain to a target domain. The CycleGAN has been utilized in several MRI harmonization studies [25, 4] and has become the foundational structure of many subsequent image-level harmonization methods.
(2) StarGAN [5] is initially introduced for multi-domain natural image synthesis task. Then, it was employed for MRI harmonization [23]. Building upon CycleGAN, the StarGAN method learns a separate latent style code for each target image and injects this code into the source image via adaptive instance normalization.
(3) ImUnity [3] is designed specifically for MRI harmonization. It employs a self-supervised variational autoencoder model, coupled with a confusion module for site/scanner-bias unlearning. In addition, it includes an optional biological preservation module to preserve clinical information, such as disease labels or diagnostic scores. In this work, we use the default form of ImUnity without this optional module, since the two datasets only consist of healthy control subjects.
In all three deep learning methods mentioned above, we typically adhere to the default setting of each competing method and make a dedicated effort to ensure other training hyperparameters align closely with the proposed DLEST framework. In DLEST, the two hyperparameters $\alpha$ and $\beta$ in Eq. 10 are empirically set as 1 and 100, respectively. For a fair comparison, we standardize the input size to be $256\times 256$ and set the latent dimension to 512, mirroring the specifications of the proposed DLEST. To train deep learning models (i.e., CycleGAN, StarGAN, ImUnity, and our DLEST), we use the OpenBHB dataset, due to its relatively larger sample size and abundant sites/settings. Following the data split in [9], $32,270$ training slices and $7,570$ validation slices are used, with manual inspection and quality control.